OpenStudy (anonymous):

Differential Equations Find the equation of the solution to dy/dx = x^3 y through the point (x,y) = (1,2)

OpenStudy (anonymous):

What I did: it's separable - 1/y dy = x^3 dx. Integrate -> ln|y| = 1/4 x^4 +c take e^both -> y = e^(1/4 x^4 +c) = ce^(1/4x^4)

OpenStudy (anonymous):

is (x,y) = (1,2) the same as: y(1) = 2 = ce^(1/4(1)^2) ?

OpenStudy (empty):

Yep!

OpenStudy (anonymous):

So c = 2e^(-1/4)

OpenStudy (empty):

Yes, exactly.

OpenStudy (anonymous):

then y = 2e^(-1/4) * e^(1/4 x^4)

OpenStudy (anonymous):

I'm getting a response that this is wrong

OpenStudy (empty):

Provided of course you're talking about the second c and not the first c you wrote, since you're reusing the letter to represent two different things, you probably should have renamed it K or something else when you took the logarithm

OpenStudy (empty):

Oh is it expecting you to simplify it more perhaps? You can combine them all into one exponent with exponent rules.

OpenStudy (anonymous):

How so?

OpenStudy (empty):

\[y=2e^{(x^4-1)/4}\] It also might be that you're not using parenthesis I'm not sure, but we can check that this does indeed satisfy the differential equation by differentiating it and plugging it in and we can also plug in x=1 and see that we get \(e^0=1\) which means we'll have y=2 like we expect.

OpenStudy (empty):

What I mean by the parenthesis bit is if you're typing in: \[y=2e^{-1/4}e^{1/4x^4}\] The computer grading mabob might be interpretting this as: \[y=2e^{-1/4}e^{1/(4x^4)}\] So you might not have to simplify it at all, but this is just myguess

OpenStudy (anonymous):

it wanted it simplified. Thanks for your help!

OpenStudy (empty):

Cool yeah ask anything anytime I'll do my best to help out, it's rare to see people who actually have tried to solve their problem and are using their mind ya know? haha xD

OpenStudy (anonymous):

Haha, thanks.